# Bayesian Projection of Refugee and Asylum Seeker Populations

### v1

@@ Abstract
Estimates of future migration patterns are a crucial input to world population projections. Forced migration, including refugee and asylum seekers, plays an important role in overall migration patterns, but is notoriously difficult to forecast. We propose a modeling pipeline based on Bayesian hierarchical time-series modeling for projecting combined refugee and asylum seeker populations by country of origin using data from the United Nations High Commissioner for Human Rights (UNHCR). Our approach is based on a conceptual model of refugee crises following growth and decline phases, separated by a peak. The growth and decline phases are modeled by logistic growth and decline through an interrupted logistic process model. We evaluate our method through a set of validation exercises that show it has good performance for forecasts at 1, 5, and 10 year horizons, and we present projections for 35 countries with ongoing refugee crises.
@@
Bayesian hierarchical model $\cdot$ Forced displacement $\cdot$ Population projections
@@ Introduction
World population projections are used for a number of reasons, including policy planning and as an input for projecting other outcomes of interest (Vanella et al., , 2020). Population projections are formed based on projections of three basic demographic variables: births, deaths, and migration (Swanson et al., , 2004). The third variable, migration, has traditionally been viewed as being harder to forecast than the first two due to inherent uncertainty in the political and economic factors that are drivers of migration (Fuchs et al., , 2021; de Valk et al., , 2022). Furthermore, forced displacement as a subset of international migration is particularly challenging to forecast because it can change rapidly; for example, the sudden outbreak of a war can lead very quickly to large numbers of refugees (Martin and Singh, , 2018). Complicating matters further is that defining who qualifies as a forced migrant is in itself a complex political and legal subject (Reed et al., , 2016).
The United Nations (UN) Population Division is a major producer of population projections through regular releases of the World Population Prospects, the most recent being from 2022 (United Nations, 2022b). The UN uses probabilistic Bayesian approaches for projecting total fertility and life expectancy (Alkema et al., , 2011; Raftery et al., , 2013). For forced migration, UN projections are currently based on an assumption that two thirds of refugees will return to their country of origin within 5 years (United Nations, 2022a). This paper is motivated by the goal of providing probabilistic projections of refugee and asylum seeker populations to serve as input for UN population projections.
@@
Our forecasting approach is based on a conceptual model of refugee crises in which crises experience growth and decline phases, separated by a peak. The growth phase corresponds to refugees leaving their country of origin due to a conflict or other crisis. The peak is a period during which the refugees stay in their country of destination. During the decline stage, the refugees return home.
@@
Trends on either side of the peak are modeled as following logistic growth (decline). We refer to this process model as an interrupted logistic model. We produce probabilistic projections using Bayesian statistical inference, drawing on the long lines of extant research applying Bayesian methods to estimate and project demographic indicators (Bijak, , 2011; Bijak and Bryant, , 2016; Raftery et al., , 2014).
@@
Other attempts have been made to apply statistical demography techniques to project forcibly displaced populations. Within the UN High Commissioner for Human Rights (UNHCR) there have been multiple efforts. The UNHCR Demographic Projection Tool applies demographic techniques to project age-specific refugee populations based on fertility and life expectency data. However, the tool does not attempt to project arrivals or departures of refugees, rather relying on the analyst to provide possible future scenarios. From another angle, the UNHCR Jetson project seeks to apply predictive modeling to forecast refugee movements (UNHCR 2024a). A predictive model of monthly arrivals of internally displaced persons by region in Somalia was developed using climate, weather, and market variables as predictors (UNHCR 2024b).
@@
Statistical and machine learning approaches have been used to predict refugee movements using indicators such as violent conflicts, market prices, and weather and climate variables (Carammia et al., , 2022; Huynh and Basu, , 2020; Singh et al., , 2019). Typically, such methods have shorter forecast horizons (on the order of weeks or months) and are targeted towards aiding in humanitarian response efforts. Similarly, a number of countries have implemented prediction systems in order to anticipate arrivals of asylum seekers (Angenendt et al., , 2023). Methods based on detecting signals in other variables are difficult to apply for long-term population projections, as doing so typically requires projecting the indicators themselves, which are often noisier than the demographic outcome being forecast. As such, our approach is based only on historical refugee and asylum seeker data. A possible downside of this approach is that our methods may be less well suited for short-term projections of the type useful for humanitarian planning in emergencies.
@@
Gravity-type models have also been applied to predict forced migration flows (Qi and Bircan, , 2023; Saldarriaga and Hua, , 2019). However, the fundamental assumptions of gravity models have been criticized as insufficient in explaining human migration patterns (Beyer et al., , 2022; Welch and Raftery, , 2022). In addition, similar to the approaches described in the previous paragraph, gravity models typically depend on complex covariates including economic, climate, and conflict variables. Projections based on gravity models therefore require projections of all covariates, which is difficult especially for long-term projections.
@@
A related methodological area of research has been the development of agent-based models (ABMs) to simulate forced migration on a variety of geographic and temporal time scales (Gray et al., , 2017; Sokolowski and Banks, , 2014; Groen, , 2016; Frydenlund et al., , 2018; Kniveton et al., , 2011). Typically ABMs simulate individual refugees according to a set of complex behavioral rules in order to form aggregate projections. However, the intensive computational requirements and unclear probabilistic properties of ABMs make them difficult to apply for longer-term forecasts, especially when statistically well-founded estimates of uncertainty are of interest.
@@ Methods
In this paper, we focus on individuals classified as refugees or asylum seekers by UNHCR based on definitions in the International Recommendations on Refugee Statistics (EGRIS, 2018). The number of refugees by country of origin were sourced from the UNHCR Refugee Population Statistics Database (UNHCR, , 2021). The UNHCR releases population statistics twice per year, with end of year statistics being released the following June. Based on definitions provided in the International Recommendations on Refugee Statistics, forcibly displaced and stateless populations are separated into five categories: refugees, asylum seekers, internally displaced persons of concern to UNHCR, other people in need of international protection, and stateless people (EGRIS, 2018). Statistics on each population type are available starting in different years depending on the host UN division.
Formally, let $t=1,\dots,T$ index years and $c=1,\dots,C$ index UN divisions (henceforth referred to generically as “countries”). The UN divisions “Unknown”, “Stateless”, “Western Sahara”, and “Palestinian” were excluded from the analysis – Western Sahara for data quality concerns (U.S. Department of State 2007), and Palestinian because those data are sourced differently than other divisions (UNHCR, , 2021). Let $R_{c,t}$ be the combined number of refugees and asylum seekers with origin country $c$ at time $t$. Let $P_{c,t}$ be the population of country $c$ at time $t$, using as source the UN World Population Prospects 2022 (United Nations, 2022b). We model refugee populations as proportions, defining $r_{c,t}=R_{c,t}/\left(R_{c,t}+P_{c,t}\right)$. The analysis dataset was formed by selecting countries with at least 10 years of observed data, at least one observed refugee proportion $r_{c,t}$ greater than 1%, and at least one observed total number of refugees $R_{c,t}$ greater than 1000.
@@
We conceptualize the number of refugees and asylum seekers from a particular country over time as arising from a succession of one or more crises, which we refer to generically as a refugee crisis although it may lead to both refugees and asylum seekers. We think of a refugee crisis as being caused by ongoing events or conditions in the country, such as war, famine, or political/economic upheaval. The refugee and asylum seeker population is expected to increase while the condition is present, before declining after the condition ends. A refugee crisis thus comprises two stages: an initial growth stage and a decline phase, separated by a peak.
Figure 1 shows illustrative data from three countries. In Mozambique, a driver of the refugee population was the Mozambican Civil War from 1977-1992. Thus, we see a sharp increase in the refugee population during the war, a peak in 1992 when the war ended, followed by a decline. Somalia appears to have experienced two crises, with the most recent crisis currently in the decline phase. Honduras appears to currently be in the growth phase of a crisis, and the timing and magnitude of the peak of the crisis is not yet known.
@@
Our modeling approach is informed by the conceptual framework for refugee crises. First, refugee and asylum seeker populations from each origin country are divided into segments corresponding to refugee crises. Next, a Bayesian hierarchical model of refugee crises is fit based on the segmented dataset. The overall pipeline is illustrated in Figure 1 using Mozambique, Somalia, and Honduras as examples.
@@
The time series for each country are first segmented into one or more refugee crises. We do so by identifying local peaks in the time series, which we take to be the peak of a corresponding refugee crisis. To identify local peaks we use a peak detection algorithm described by Palshikar et al., (2009) and implemented in the scorepeak package (Ochi, , 2019). Once the peaks are identified, we identify the extent of the refugee crisis by looking for troughs before and after the peak. A trough is defined as the minimum proportion observed between two peaks (or between a peak and the beginning or end of the time series). The start of a crisis is defined as the first year after a trough (or the beginning of the time-series) in which the refugee proportion $r_{c,t}$ is greater than 0.025%. If such a year does not exist, the start is considered to be the year before the peak.
Each crisis is next classified as ended or ongoing. A crisis is classified as ended if it is either followed by another crisis, or if the most recent observed refugee proportion $r_{c,t}$ is less than 0.025%. For an ended crisis, the end time point is set to the year in which the last observation was greater than 0.025%. All other crises are classified as ongoing. Ongoing crisis are further divided into being in the growth or decline phase. Countries that have a detected local peak on or after 2015, or that have a refugee proportion in 2021 greater than 80% of the most recent peak, are classified as being in the growth phase. All other ongoing crises are classified as being in the decline phase.
Figure 1 illustrates the detected crises for Mozambique, Somalia, and Honduras. Two crises are detected in Mozambique, corresponding roughly to the Mozambican War of Independence and the Mozambican Civil War. The second crisis in Mozambique is classified as having ended, as all observations after 1995 are under 0.025%. Somalia also has two crises detected, corresponding to different phases of the Somali Civil War. The second crisis in Somalia is classified as ongoing and in the decline phase, as the most recent observations are all over 0.025%. Finally, an ongoing crisis is also detected in Honduras, but is classified as being in the growth phase. Table 1 shows the classifications of the final crisis in each country in the analysis dataset.
Table 1: Countries in the analysis dataset grouped by classification of their most recent crisis.
Ended crises
@@
Ongoing crises, decline phase
@@
Ongoing crises, growth phase
@@
Formally, let $m=1,\dots,M$ index the detected refugee crises from the segmentation step, and denote by $c[m]$ the country corresponding to the $m$th crisis. Let $t^{\mathrm{start}}_{m}$, $t^{\mathrm{end}}_{m}$ be the start and end time points of the $m$th crisis, respectively. Let $t^{\mathrm{peak}}_{m}$ be the time point of the peak of crisis $m$.
@@
Let $\mu_{c,t}$ denote the modeled refugee proportion in country $c$ at time $t$. The process model describes how $\mu_{c,t}$ is expected to evolve over time (Susmann et al., , 2022). The core of our modeling approach is to assume that refugee populations follow logistic growth and decline trends during the growth and decline phases of a refugee crisis, respectively. The rate of change of $\mu_{c[m],t}$ during the growth and decline phases is modeled by the following logistic rate function:
@@
where $\omega\in\mathbb{R}$ is a rate parameter and $\lambda\in(0,1)$ is an asymptote parameter. During the growth phase, a crisis-specific rate parameter $\omega_{m}^{(1)}>0$ is applied. During the decline phase a rate parameter $\omega_{m}^{(2)}<0$ is used. Both growth and decline phases share an asymptote parameter $\lambda_{m}$. Formally, the expected rate of change for crisis $m$ at time $t$ is given by
@@
We refer to the process model as an interrupted logistic process model because the logistic growth during the growth phase is interrupted at time $t^{\mathrm{peak}}_{m}$. Figure 2 illustrates the parameters of the interrupted logistic model for a single crisis.
@@
Crisis lengths and maxima
We model the length of the growth phase and the refugee proportion observed at the peak for use in projecting crises that are still in the growth phase. The length of the growth phase is assumed to be exponentially distributed with rate parameter $\psi>0$:
$t^{\mathrm{peak}}_{m}-t^{\mathrm{start}}_{m}\sim\mathrm{Exp}(\psi).$ (5)
The exponential model was chosen based on empirical inspection of the distribution of growth phase lengths (Appendix Figure 6). The proportion at the peak is assumed to follow a logit-normal distribution around a mean parameter $\rho$:
@@
where $\mathrm{logit}(x)=\log(x/(1-x))$.
@@
Hierarchical priors are used to share information between crises. The upper asymptote of each crisis is assumed to follow a logit-normal distribution around an overall mean:
@@
$\sigma_{c}^{(1)}$ $\sim N(\mu_{\sigma^{(1)}},\sigma_{\sigma^{(1)}}^{2}),$ (10)
$\sigma_{c}^{(2)}$ $\sim N(\mu_{\sigma^{(2)}},\sigma_{\sigma^{(2)}}^{2}).$ (11)
@@
Projections are produced for all ongoing crises. Suppose we have at our disposal $k=1,\dots,K$ posterior draws from the joint posterior distribution of the model parameters conditional on all the observed data. We will write $\alpha^{(k)}$ to denote the $k$th posterior draw of a parameter $\alpha$.
@@
The projection method is different depending on whether the ongoing crisis was classified as being in the decline or growth phase. For a crisis $m$ that is ongoing and in the decline phase, the timing and level of the peak is taken as fixed. Thus, for all $t>t^{\mathrm{last}}_{m}$, the refugee proportion is projected by recursive application of the following equation for each posterior draw $k=1,\dots,K$:
$\mu_{c[m],t}^{(k)}\sim N\left(\mu_{c[m],t-1}^{(k)}+f\left(\mu_{c[m],t-1}^{(k)},\omega_{m}^{(2)(k)},\lambda_{m}^{(k)}\right),\left(\sigma^{(k)}_{c,t}\right)^{2}\right),$ (13)
@@
and where we set the initial value $\mu_{c[m],t^{\mathrm{last}}_{m}}^{(k)}=r_{c[m],t^{\mathrm{last}}_{m}}$. If, for any time $t$, $\mu_{c[m],t}^{(k)}<0$, then we set $\mu_{c[m],t}^{(k)}=0.001$ before continuing recursive application of the formula.
@@
For an ongoing crisis $m$ in the growth phase, the timing and level of the peak has not been observed, and we integrate over this uncertainty. Suppose the peak falls in year $t^{\mathrm{last}}_{m}+\delta$, where $\delta\in\mathbb{N}^{+}$ (in practice we take $\delta\in\{0,1,\dots,15\})$. Conditional on $\delta$, projections are generated by recursive application of the equation
@@
$\mu_{c[m],t+1,\delta}^{(k)}\sim\begin{cases}N\left(\mu_{c[m],t,\delta}^{(k)}+f\left(\mu_{c[m],t,\delta}^{(k)},\omega_{m}^{(1)(k)}\right),\lambda_{m}^{(k)},(\sigma^{(k)}_{c,t})^{2}\right),&t\leq t^{\mathrm{last}}_{m}+\delta,\\ N\left(\mu_{c[m],t,\delta}^{(k)}+f\left(\mu_{c[m],t,\delta}^{(k)},\omega_{m}^{(2)(k)}\right),\lambda_{m}^{(k)},(\sigma^{(k)}_{c,t})^{2}\right),&t>t^{\mathrm{last}}_{m}+\delta.\end{cases}$ (14)
@@
As before, if for any $t$ $\mu_{c[m],t,\delta}^{(k)}<0$, then we set $\mu_{c[m],t,\delta}^{(k)}=0.001$. Next, the probability of $\delta$ conditional on the implied crisis length and the refugee proportion at the peak is calculated conditional on the crisis length parameters $\psi$ and peak proportion parameters $\rho$ and $\sigma_{\rho}^{2}$ previously estimated.
$p_{\delta}\propto d_{\mathrm{Exp}}(\delta\mid\psi^{(k)})d_{\mathrm{N}}(\mathrm{logit}(\mu_{c[m],t^{\mathrm{last}}_{m}+\delta,\delta}^{(k)})\mid\rho,\sigma^{2}_{\rho}),$ (15)

### v2

# Bayesian Projection of Extant Refugee and Asylum Seeker Populations

@@ Abstract
Estimates of future migration patterns are of broad interest in demography. Forced migration, including refugee and asylum seekers, plays an important role in overall migration patterns, but is notoriously difficult to forecast. Focusing on refugees and asylum seekers, we propose a modeling pipeline based on Bayesian hierarchical time-series modeling for projecting refugee population official statistics by country of origin using data from the United Nations High Commissioner for Refugees (UNHCR). Our approach is based on a conceptual model of refugee and asylum seeker populations following growth and decline phases, separated by a peak. The growth and decline phases are modeled by logistic growth and decline through an interrupted logistic process model. We evaluate our method through a set of validation exercises that show it has good performance for forecasts at 1, 5, and 10 year horizons, and we present projections for 35 countries of origin of large refugee and asylum seeker population.
@@
Bayesian hierarchical model $\cdot$ Refugees $\cdot$ Population projections
@@ Introduction
The study of refugee and asylum seeker populations is a large and evolving research area in demography (Arar and Fitzgerald, , 2023; Hugo et al., , 2018). Migration has traditionally been viewed as difficult to forecast due to inherent uncertainty in political and economic factors that are drivers of migration (Fuchs et al., , 2021; de Valk et al., , 2022), and refugee and asylum seeker populations as a subset of international migration are particularly challenging to forecast. For example, the sudden outbreak of a war can lead very quickly to large numbers of refugees (Martin and Singh, , 2018). Complicating matters further is that defining who qualifies as a refugee or asylum seeker is in itself a complex political and legal subject (Reed et al., , 2016).
@@
We propose an approach to projecting refugee and asylum seeker statistics by country of origin. Importantly, we focus on producing projections for countries with extant large reported populations of refugees and asylum seekers; we do not attempt to forecast which countries will begin to be new origins of refugees and asylum seekers. Our forecasting approach is based on a conceptual model of refugee and asylum seeker populations (stocks) in which these populations experience growth and decline phases, separated by a peak. The growth phase corresponds to refugees and asylum seekers leaving their country of origin due to a conflict or other reason. The peak is a period during which the refugees and asylum seeker stay in their country of destination, or move to a alternate host country. During the decline stage, the refugees and asylum seekers return home or lose protected status for other reasons, such naturalization in their host country.
@@
Trends on either side of the peak are modeled as following logistic growth before the peak and decline after it. We refer to this process model as an interrupted logistic model. We produce probabilistic projections of combined refugee and asylum seeker stocks using Bayesian statistical inference, drawing on the long line of previous research applying Bayesian methods to estimate and project demographic indicators (Bijak, , 2011; Bijak and Bryant, , 2016; Azose and Raftery, , 2015; Azose et al., , 2016).
The first step in modeling refugee and asylum seeker counts is to precisely define what is meant by these two terms. Other demographic indicators of interest, such as fertility or mortality rates, are defined by reference to immutable and universally recognized events. In this sense, the definition of what qualifies as a birth or a death is not controversial. The same definitional clarity is not manifest when it comes to tracking refugee and asylum seeker populations, as deciding who qualifies as a refugee or asylum seeker is in itself a complex political and legal subject that is continually evolving. In this work, we choose to define refugees and asylum seekers following definitions in the International Recommendations on Refugee Statistics (EGRIS, 2018). This choice is in part practical, in that the UN High Commissioner for Refugees (UNHCR) releases official statistics of refugee and asylum seeker populations based on these definitions (UNHCR, , 2021). We emphasize that our projections of refugee and asylum seeker populations are therefore to be solely understood as projections of refugee and asylum seeker populations as those populations are defined by the EGRIS recommendations and reflected in UNHCR statistics.
Long-term forecasts of refugee populations have been produced as inputs to world population projections; in particular, the United Nations (UN) Population Division is a major producer of population projections through regular releases of the World Population Prospects. The UN uses probabilistic Bayesian approaches for projecting total fertility and life expectancy (Alkema et al., , 2011; Raftery et al., , 2013, 2014). For refugees, UN projections are currently based on an assumption that two thirds of refugees will return to their country of origin within 5 years (United Nations, 2022a). We adopt this simple method as a benchmark against which to compare our proposed approach.
@@
Other attempts have been made to apply techniques from statistical demography to project forcibly displaced populations. Within the UN High Commission for Refugees (UNHCR) there have been multiple such efforts. The UNHCR Demographic Projection Tool applies demographic methods to project age-specific refugee populations based on fertility and life expectency data. However, the tool does not attempt to project arrivals or departures of refugees, rather relying on the analyst to provide possible future scenarios. From another angle, the UNHCR Jetson project seeks to apply predictive modeling to forecast refugee movements (UNHCR 2024a). A predictive model of monthly arrivals of internally displaced persons by region in Somalia was developed using climate, weather, and market variables as predictors (UNHCR 2024b).
@@
Statistical and machine learning approaches have been used to predict refugee and asylum seeker movements using indicators such as violent conflicts, market prices, and weather and climate variables (Carammia et al., , 2022; Huynh and Basu, , 2020; Singh et al., , 2019). Typically, such methods have shorter forecast horizons (on the order of weeks or months) and are targeted towards aiding in humanitarian response efforts. Similarly, a number of countries have implemented prediction systems in order to anticipate arrivals of asylum seekers (Angenendt et al., , 2023). Methods based on detecting signals in other variables are difficult to apply for medium- to long-term population projections, as doing so typically requires projecting the indicators themselves, which are often noisier than the demographic outcome being forecast. As such, our approach is based only on historical refugee data. A possible downside of this approach is that our methods may be less well suited for short-term projections of the type useful for humanitarian planning in emergencies.
@@
Gravity-type models have also been applied to predict forced migration flows (Qi and Bircan, , 2023; Saldarriaga and Hua, , 2019). However, the fundamental assumptions of gravity models have been criticized as insufficient in explaining human migration patterns (Beyer et al., , 2022; Welch and Raftery, , 2022). In addition, similarly to the approaches described in the previous paragraph, gravity models typically depend on complex covariates including economic, climate, and conflict variables. Projections based on gravity models therefore require projections of all covariates, which is difficult especially for longer-term projections.
@@
A related methodological area of research has been the development of agent-based models (ABMs) to simulate forced migration on a variety of geographic and temporal time scales (Gray et al., , 2017; Sokolowski and Banks, , 2014; Groen, , 2016; Frydenlund et al., , 2018; Kniveton et al., , 2011). Typically ABMs simulate individual refugees according to a set of complex behavioral rules. ABMs may be developed for reasons other than forecasting, such as simulating the effect of interventions on outcomes of interest. The intensive computational requirements and unclear probabilistic properties of ABMs complicates their application to producing longer-term forecasts, especially when statistically well-founded estimates of uncertainty are of interest. Recent work on ABM simulations of migration incorporating uncertainty via Bayesian principles provides a potential path forward (Bijak et al., , 2021).
@@ Methods
We focus on individuals classified as refugees and asylum seekers by UNHCR based on definitions in the International Recommendations on Refugee Statistics (EGRIS, 2018). The number of refugees and asylum seekers by country of origin were sourced from the UNHCR Refugee Population Statistics Database (UNHCR, , 2021). Refugees are those “persons who have current refugee status, granted either before arrival or upon arrival in the receiving country” (EGRIS, 2018; see within item 92.3 for extended definition, and Chapter 2 for legal discussion of refugee status). Asylum seekers are defined as “Persons who have filed an application for asylum in a country other than their own and whose claims have not yet been determined. These include those filing primary applications or subsequent applications following an appeal. The date on which the application for asylum is filed marks their entry into the status of asylum seeker. They remain in the status of asylum seeker until their application is considered and adjudicated.” (92.2 EGRIS, 2018)
The UNHCR releases population statistics twice per year, with end of year statistics being released the following June. Based on definitions provided in the International Recommendations on Refugee Statistics, forcibly displaced and stateless populations are separated into five categories: refugees, asylum seekers, internally displaced persons of concern to UNHCR, other people in need of international protection, and stateless people (EGRIS, 2018). Statistics on each population type are available starting in different years depending on the host UN division.
Formally, let $t=1,\dots,T$ index years and $c=1,\dots,C$ index UN divisions (henceforth referred to generically as “countries”). The UN divisions “Unknown”, “Stateless”, “Western Sahara”, and “Palestinian” were excluded from the analysis – Western Sahara for data quality concerns (U.S. Department of State 2007), and Palestinian because those data are sourced differently than other divisions (UNHCR, , 2021). Let $R_{c,t}$ be the combined number of refugees and asylum seekers with origin country $c$ at time $t$. Let $P_{c,t}$ be the population of country $c$ at time $t$, using as source the UN World Population Prospects 2022 (United Nations, 2022b). We model refugee and asylum seeker populations as proportions, defining $r_{c,t}=R_{c,t}/\penalty 50\left(R_{c,t}+P_{c,t}\right)$. The analysis dataset was formed by selecting countries with at least 10 years of observed data, at least one observed proportion $r_{c,t}$ greater than 1%, and at least one observed population size $R_{c,t}$ greater than 1000.
@@
We conceptualize the number of refugees and asylum seekers from a particular country over time as arising from a succession of one or more periods characterized by growth and subsequent decline which we refer to generically as a growth/decline period or, succinctly, as an episode. We think of an episode as being caused by ongoing events or conditions in the country, such as war, famine, or political/economic upheaval. The refugee and asylum seeker population is expected to increase while the condition is present, before declining after the condition ends. An episode thus comprises two stages: an initial growth stage and a decline phase, separated by a peak. We note that a notable limitation of our conceptual framework is that it may not well capture long-term refugee populations in which population dynamics are dominated by traditional demographic forces.
Figure 1 shows illustrative data from three countries. In Mozambique, a driver of the refugee and asylum seeker population was the Mozambican Civil War from 1977-1992. Thus, we see a sharp increase in the refugee and asylum seeker population during the war, a peak in 1992 when the war ended, followed by a decline. Somalia appears to have experienced two episodes, with the most recent episode currently in the decline phase. Honduras appears to currently be in the growth phase of a episode, and the timing and magnitude of the peak of the episode is not yet known.
@@
Our modeling approach is informed by the conceptual framework for episodes. First, refugee and asylum seeker populations from each origin country are divided into growth/decline periods. Next, a Bayesian hierarchical model is fit based on the segmented dataset. The overall pipeline is illustrated in Figure 1 using Mozambique, Somalia, and Honduras as examples.
@@
The time series for each country is first segmented into one or more episodes of growth and decline. We do this by identifying local peaks in the time series, which we take to be the peak of a corresponding episode. To identify local peaks we use a peak detection algorithm described by Palshikar et al., (2009) and implemented in the scorepeak package (Ochi, , 2019). Once the peaks are identified, we identify the extent of the episode by looking for troughs before and after the peak. A trough is defined as the minimum proportion observed between two peaks (or between a peak and the beginning or end of the time series). The start of a episode is defined as the first year after a trough (or the beginning of the time-series) in which the refugee and asylum seeker proportion $r_{c,t}$ is greater than 0.025%. If such a year does not exist, the start is considered to be the year before the peak.
Each episode is next classified as ended or ongoing. An episode is classified as ended if it is either followed by another episode, or if the most recent observed proportion $r_{c,t}$ is less than 0.025%. For an ended episode, the end time point is set to the year in which the last observation was greater than 0.025%. All other episodes are classified as ongoing. Ongoing episodes are further divided into being in the growth or decline phase. Countries that have a detected local peak on or after 2015, or that have a proportion in 2021 greater than 80% of the most recent peak, are classified as being in the growth phase. All other ongoing episodes are classified as being in the decline phase.
Figure 1 illustrates the detected episodes for Mozambique, Somalia, and Honduras. Two episodes are detected in Mozambique, corresponding roughly to the Mozambican War of Independence and the Mozambican Civil War. The second episode in Mozambique is classified as having ended, as all observations after 1995 are under 0.025%. Somalia also has two episodes detected, corresponding to different phases of the Somali Civil War. The second episode in Somalia is classified as ongoing and in the decline phase, as the most recent observations are all over 0.025%. Finally, an ongoing episode is also detected in Honduras, but is classified as being in the growth phase. Table 1 shows the classifications of the final episode in each country in the analysis dataset.
Table 1: Countries in the analysis dataset grouped by classification of their most recent episode. Country names are presented as in the data source.
Ended episodes
@@
Ongoing episodes, decline phase
@@
Ongoing episodes, growth phase
@@
Formally, let $m=1,\dots,M$ index the detected episodes from the segmentation step, and denote by $c[m]$ the country corresponding to the $m$th episode. Let $t^{\mathrm{start}}_{m}$, $t^{\mathrm{end}}_{m}$ be the start and end time points of the $m$th episode, respectively. Let $t^{\mathrm{peak}}_{m}$ be the time point of the peak of episode $m$.
@@
Let $\mu_{c,t}$ denote the modeled refugee proportion in country $c$ at time $t$. The process model describes how $\mu_{c,t}$ is expected to evolve over time (Susmann et al., , 2022). The core of our modeling approach is to assume that refugee and asylum seeker populations follow logistic growth and decline trends during the growth and decline phases of a episode, respectively. The rate of change of $\mu_{c[m],t}$ during the growth and decline phases is modeled by the following logistic rate function:
@@
where $\omega\in\mathbb{R}$ is a rate parameter and $\lambda\in(0,1)$ is an asymptote parameter. During the growth phase, an episode-specific rate parameter $\omega_{m}^{(1)}>0$ is applied. During the decline phase a rate parameter $\omega_{m}^{(2)}<0$ is used. Both growth and decline phases share an asymptote parameter $\lambda_{m}$. Formally, the expected rate of change for episode $m$ at time $t$ is given by
@@
We refer to the process model as an interrupted logistic process model because the logistic growth during the growth phase is interrupted at time $t^{\mathrm{peak}}_{m}$. Figure 2 illustrates the parameters of the interrupted logistic model for a single episode.
@@
Episode lengths and maxima
We model the length of the growth phase and the refugee and asylum seeker proportion observed at the peak for use in projecting episodes that are still in the growth phase. The length of the growth phase is assumed to be exponentially distributed with rate parameter $\psi>0$:
$t^{\mathrm{peak}}_{m}-t^{\mathrm{start}}_{m}\sim\mathrm{Exp}(\psi),$ (5)
with prior $\psi\sim\mathrm{InvGamma}(0.1,0.1)$, an inverse Gamma distribution. The exponential model was chosen based on empirical inspection of the distribution of growth phase lengths (Appendix Figure A1); we note also the natural connection between the exponential distribution and the distribution of waiting times. The proportion at the peak is assumed to follow a logit-normal distribution around a mean parameter $\rho$:
@@
where $\mathrm{logit}(x)=\log(x/(1-x))$. The following priors are assigned to $\rho$ and $\sigma_{\rho}$:
$\rho$ $\sim N(0,3),$ (7)
$\sigma_{\rho}$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (8)
@@
Hierarchical priors are used to share information between episodes. The upper asymptote of each episode is assumed to follow a logit-normal distribution around an overall mean:
@@
$\log\sigma_{c}^{(1)}$ $\sim N(\mu_{\sigma^{(1)}},\sigma_{\sigma^{(1)}}^{2}),$ (13)
$\log\sigma_{c}^{(2)}$ $\sim N(\mu_{\sigma^{(2)}},\sigma_{\sigma^{(2)}}^{2}).$ (14)
The priors used for the hierarchical models are as follows, for the growth and decline rates:
$\mu_{\omega^{(1)}}$ $\sim N(0,3),$ (16)
$\mu_{\omega^{(2)}}$ $\sim N(0,3),$ (17)
$\sigma_{\omega^{(1)}}$ $\sim\mathrm{InvGamma}(0.1,0.1),$ (18)
$\sigma_{\omega^{(2)}}$ $\sim\mathrm{InvGamma}(0.1,0.1),$ (19)
and for the growth and decline noise parameters:
$\mu_{\sigma^{(1)}}$ $\sim N(-7,4),$ (20)
$\mu_{\sigma^{(2)}}$ $\sim N(-7,4),$ (21)
$\sigma_{\sigma^{(1)}}$ $\sim\mathrm{InvGamma}(0.1,0.1)$ (22)
$\sigma_{\sigma^{(2)}}$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (23)
@@
Projections are produced for all ongoing episodes. Suppose we have at our disposal $k=1,\dots,K$ posterior draws from the joint posterior distribution of the model parameters conditional on all the observed data. We will append an additional suffix to denote the $k$th posterior draw, as in $\alpha_{k}$ to denote the $k$th posterior draw of the parameter $\alpha$.
@@
The projection method is different depending on whether the ongoing episode was classified as being in the decline or growth phase. For a episode $m$ that is ongoing and in the decline phase, the timing and level of the peak is taken as fixed. Thus, for all $t>t^{\mathrm{last}}_{m}$, the refugee and asylum seeker proportion is projected by recursive application of the following equation for each posterior draw $k=1,\dots,K$:
$\mu_{c[m],t,k}\sim N\left(\mu_{c[m],t-1,k}+f\left(\mu_{c[m],t-1,k},\omega_{m,k}^{(2)},\lambda_{m,k}\right),\sigma^{2}_{c,t,k}\right),$ (24)
@@
and where we set the initial value $\mu_{c[m],t^{\mathrm{last}}_{m},k}=r_{c[m],t^{\mathrm{last}}_{m}}$. If, for any time $t$, $\mu_{c[m],t,k}<0$, then we set $\mu_{c[m],t,k}=0.001$ before continuing recursive application of the formula.
@@
For an ongoing episode $m$ in the growth phase, the timing and level of the peak has not been observed, and we integrate over this uncertainty. Suppose the peak falls in year $t^{\mathrm{last}}_{m}+\delta$, where $\delta\in\mathbb{N}^{+}$ (in practice we take $\delta\in\{0,1,\dots,15\})$. Conditional on $\delta$, projections are generated by recursive application of the equation
@@
$\mu_{c[m],t+1,\delta,k}\sim\begin{cases}N\left(\mu_{c[m],t,\delta,k}+f\left(\mu_{c[m],t,\delta,k},\omega_{m,k}^{(1)}\right),\lambda_{m,k},\sigma^{2}_{c,t,k}\right),&t\leq t^{\mathrm{last}}_{m}+\delta,\\ N\left(\mu_{c[m],t,\delta,k}+f\left(\mu_{c[m],t,\delta,k},\omega_{m,k}^{(2)}\right),\lambda_{m,k},\sigma^{2}_{c,t,k}\right),&t>t^{\mathrm{last}}_{m}+\delta.\end{cases}$ (25)
@@
As before, if for any $t$ $\mu_{c[m],t,\delta,k}<0$, then we set $\mu_{c[m],t,\delta,k}=0.001$. Next, the probability of $\delta$ conditional on the implied episode length and the refugee and asylum seeker proportion at the peak is calculated conditional on the episode length parameters $\psi$ and peak proportion parameters $\rho$ and $\sigma_{\rho}^{2}$ previously estimated.
$p_{\delta}\propto f_{\mathrm{Exp}}(\delta\mid\psi_{k})f_{\mathrm{N}}(\mathrm{logit}(\mu_{c[m],t^{\mathrm{last}}_{m}+\delta,\delta,k})\mid\rho,\sigma^{2}_{\rho}),$ (26)